# Token Pruning in Multimodal Large Language Models: Are We Solving the Right Problem?

### v1

@@ Abstract
Zichen Wen1,2111Equal Contribution., Yifeng Gao1111Equal Contribution., Weijia Li3,2, Conghui He2222Corresponding authors., Linfeng Zhang1222Corresponding authors. 1Shanghai Jiao Tong University 2Shanghai AI Laboratory 3Sun Yat-sen University zichen.wen@outlook.com, heconghui@pjlab.org.cn, zhanglinfeng@sjtu.edu.cn
@@ 1 Introduction
To solve this problem, abundant efforts have been made in token pruning (Chen et al., 2024; Zhang et al., 2024b; Liu et al., 2024d), which aims to reduce the number of input tokens in MLLMs. Usually, token pruning methods first introduce a carefully-designed criterion to measure the importance of a vision token, and then prune the redundant tokens, or merge the redundant tokens into fewer tokens. As a result, the following computation of the pruned tokens or the merged tokens can be removed or reduced, bringing efficiency in both computation and storage. For instance, some recent works show that more than 70% tokens can be pruned with tolerant loss in accuracy (Chen et al., 2024). Most attractively, thanks to the natural ability of MLLMs to process tokens in different lengths, token pruning can be applied to most existing MLLMs with no need for additional training, and thus attracts great attention from both academic researchers and industrial developers.
@@
However, despite the popularity of token pruning, numerous foundational questions have long been overlooked and remain largely unexplored, giving rise to several surprising phenomena. For instance, Figure 1 demonstrates the comparison between two classical token pruning methods including FastV (Chen et al., 2024) and SparseVLM (Zhang et al., 2024b), and two naive baselines, including random token selection and direct average pooling on tokens. Surprisingly, the two baselines outperform the two well-designed token pruning methods in most benchmarks by a clear margin. This counterintuitive phenomenon may demonstrate that the current understanding of so-called important tokens is far away from the truth. Unfortunately, most recent works just focus on purchasing higher performance, while ignoring these questions, which may hinder the long-term development of token pruning.
@@ 2.1 Multimodal Large Language Models
The remarkable success of large language models (LLMs) (Radford et al., 2019; Brown et al., 2020) has spurred a growing trend of extending their advanced reasoning capabilities to multi-modal tasks, leading to the development of vision-language models (VLMs) (Huang et al., 2023; Driess et al., 2023; Liu et al., 2024c; Bai et al., 2023). These VLMs typically consist of a visual encoder (Radford et al., 2021) that serializes input image representations and an LLM responsible for text generation. To enable the LLM to process visual inputs, an alignment module is employed to bridge the gap between visual and textual modalities. This module can take various forms, such as a simple linear layer, an MLP projector, or a more complex query-based network. While this integration allows the LLM to gain visual perception, it also introduces significant computational challenges due to the long sequences of visual tokens.
@@ 2.2 Visual Token Compression
Visual tokens are often significantly more numerous than text tokens, with higher spatial redundancy and lower information density. To address this issue, various methods have been proposed for reducing visual token counts in vision language models. For instance, some approaches modify model components, such as using context tokens in Q-Former (Li et al., 2024) or applying adaptive pooling at the patch level, but these typically require additional training and increase computational costs. Other techniques, like Token Merging (ToMe) (Bolya et al., 2023) and FastV (Chen et al., 2024), focus on reducing tokens without retraining by merging tokens or selecting important ones based on attention scores. SparseVLM Zhang et al. (2024b) incorporates text guidance through cross-modal attention to refine token selection. However, these methods often overlook hardware acceleration compatibility and fail to account for token duplication alongside token importance. Furthermore, while token pruning has been extensively explored in natural language processing and computer vision to improve inference efficiency, its application to VLMs remains under-explored. Existing pruning strategies, such as those in FastV and SparseVLM, rely on text-visual attention within large language models (LLMs) to evaluate token importance, which may not align well with actual visual token relevance.
@@ 3 Benchmarking
We begin by presenting the datasets, models, and pruning methods included in our study, along with the rationale behind selection. Next, we outline the experimental setup and provide guidance on interpreting the results reported in our study. Finally, we analyze the findings, emphasizing notable patterns and offering insights that may inform future research in this area.
@@ 3.1 Models
We selected several representative open-source MLLMs, including LLaVA-1.5-7B Liu et al. (2024a), LLaVA-Next-7B (Liu et al., 2024b), and Qwen2-VL (Wang et al., 2024) series (7B-Instruct and 72B-Instruct). LLaVA-1.5-7B integrates CLIP and LLaMA for vision-language alignment via end-to-end training, employing MLP connectors to fuse visual-text features for multimodal reasoning. LLaVA-Next-7B enhances data efficiency and inference robustness with dynamic resolution and hierarchical feature integration, improving fine-grained visual understanding. Qwen2-VL series excel in high-resolution input processing and instruction-following, supporting complex tasks like document analysis and cross-modal in-context learning through unified vision-language representations.
@@ 3.3 Token Pruning Method
To rigorously evaluate the properties of visual token pruning, we select three representative and high-performing methods: FastV (Chen et al., 2024), SparseVLM Zhang et al. (2024b), and MustDrop (Liu et al., 2024d). FastV Chen et al. (2024) optimizes computational efficiency by learning adaptive attention patterns in early layers and pruning low-attention visual tokens post-layer 2 of LLMs, effectively reducing redundancy. SparseVLM (Zhang et al., 2024b) introduces a text-guided, training-free pruning mechanism that leverages self-attention matrices between text and visual tokens to assess importance. It maximizes sparsity while preserving semantically relevant tokens without additional parameters or fine-tuning. MustDrop (Liu et al., 2024d) addresses token redundancy across the entire model lifecycle. It merges spatially similar tokens during vision encoding, employs text-guided dual-attention filtering in prefilling, and implements output-aware KV cache compression during decoding. This multi-stage approach ensures balanced retention of critical tokens while enhancing inference efficiency. These methods exemplify diverse strategies for token pruning, spanning adaptive attention, text-guided sparsity, and lifecycle-aware optimization.
@@ 4 Token Pruning Revisited: Are Simple Methods Better?
When considering token pruning in multimodal large language models, two very basic methods naturally come to mind: random token pruning (hereafter referred to as Random) and token pooling (hereafter referred to as Pooling). Comparison with these two simple baselines is reliable evidence to demonstrate the significance of a well-designed token pruning method, yet has been ignored by most previous works. To address this gap, we investigated these two simple approaches in detail. Specifically, we conducted experiments on multiple widely-used benchmarks under pruning ratios of 75% and 87.5%, comparing Random and Pooling [^1] In the experiments, Pooling specifically refers to applying a pooling operation to the visual tokens at the second layer of the language model. Please refer to the specific implementation in Algorithm 4. with several recent token pruning methods (e.g., FastV and SparseVLM). As shown in Table 1, surprisingly, Random and Pooling outperformed carefully designed methods on nearly $2/3$ benchmarks. These surprising results shocked us since its inferior performance compared with random selection may demonstrate that we are on the wrong road toward the ideal token pruning methods.
@@ 4.1 Token Distribution: Spatial Uniformity Outperforms Position Bias
Figure 3: Sparse Visualization of Vanilla FastV and Window FastV with 75% Retained Visual Tokens.
@@ 5 Language in Visual Token Pruning: When and Why Does Language Matter?
Token pruning methods for multimodal models can be broadly categorized into two types: those guided by textual information (e.g., FastV Chen et al. (2024), SparseVLM Zhang et al. (2024b), MustDrop Liu et al. (2024d)) and those that rely solely on visual information (e.g., FasterVLM Zhang et al. (2024a)). While both approaches achieve comparable performance on common benchmarks, however, we hypothesize: Could it be that the importance of language information is not evident simply because there has been a lack of testing on tasks where language information is especially critical? To validate our hypothesis, we select a typical scenario: Visual Haystack.
@@ 5.1 Visual Token Pruning in Strongly Text-Guided Tasks
However, there are also recent works methods (Zhang et al., 2024a; Liu et al., 2025a) that perform pruning solely in ViT without textual information and reports better performance than FastV and SparseVLM in common VQA benchmarks.
@@ 6.1 Redundancy Criteria
Through the lens of mutual information (Latham and Roudi, 2009), we formulate this as maximizing information preservation between original tokens $X$ and retained tokens $X^{\prime}$:
$\max_{\mathcal{P}}I(X;X^{\prime})=\mathcal{H}(X)-\mathcal{H}(X|X^{\prime}),$ (1)
where $\mathcal{P}$ denotes the pruning operator. This ensures the $X^{\prime}$ retains maximal dependence on $X$ under length constraint $\|X^{\prime}\|=\|X\|-\Delta L$. The formulation directly connects to the compression phase of the information bottleneck principle (Tishby et al., 2000), where $\mathcal{P}^{*}$ solves:
$\mathcal{P}^{*}=\operatorname*{arg\,min}_{\mathcal{P}}\|X^{\prime}\|\quad\text{s.t.}\ I(X;X^{\prime})\geq\gamma,$ (2)
@@ 6.2 Importance Criteria
In contrast, this task-oriented criterion explicitly considers the target output $Y$. The goal shifts to preserving tokens critical for prediction accuracy, formalized through predictive sufficiency:
$I(X^{\prime};Y)\geq I(X;Y)-\epsilon,$ (3)
@@
$\underbrace{I(X;Y)}_{\text{Original}}=\underbrace{I(X^{\prime};Y)}_{\text{Pruned}}+\underbrace{I(X\setminus X^{\prime};Y|X^{\prime})}_{\text{Discarded}}.$ (4)
The bound $I(X\setminus X^{\prime};Y|X^{\prime})\leq\epsilon$ implies that discarded tokens provide negligible additional information about $Y$ when conditioned on retained tokens. This captures the essence of importance - truly critical tokens contain non-decomposable predictive information.
@@
$\mathcal{R}(\beta)=\max_{X^{\prime}}\left[I(X^{\prime};Y)-\beta^{-1}I(X;X^{\prime})\right],$ (5)
@@ 6.3 Empirical Validation of Adaptive Criteria Balancing
$\alpha\cdot\underbrace{I(x_{i};Y|x_{\setminus i})}_{\text{Predictive Criticality}}+(1-\alpha)\cdot\underbrace{[1-I(x_{i};X_{\setminus i})]}_{\text{Pattern Uniqueness}}.$ (6)
Here $I(x_{i};Y|x_{\setminus i})$ measures a token’s unique predictive value, while $1-I(x_{i};X_{\setminus i})$ quantifies its pattern distinctiveness.
@@
  • Perception-Dominant Tasks (MME, POPE) achieve peak performance at $\alpha=0.1$ and $0.0$, respectively, favoring redundancy-first pruning to maintain structural integrity ($\uparrow I(X;X^{\prime})$).
@@ Appendix A Future Works
In this work, we have conducted an in-depth exploration of a series of issues related to token pruning. These include the position bias problem inherent in methods based on attention scores, the guiding role of linguistic information during token pruning, the importance and redundancy of tokens, as well as certain limitations in the evaluation of token pruning methods. Looking ahead, we plan to further expand the scope of our research by considering whether token pruning or token merging should be prioritized in the context of token reduction. Additionally, we aim to evaluate and analyze various token reduction methods on more challenging OCR benchmarks, particularly datasets featuring rich-text OCR images. This future work will not only deepen our understanding of token reduction strategies but also provide valuable insights into their practical applications in complex scenarios.

### v2

@@ Abstract
Zichen Wen1,2,3111Equal Contribution.  Yifeng Gao1,2111Equal Contribution.  Weijia Li4,3  Conghui He3222Corresponding authors.  Linfeng Zhang1,$2$Corresponding authors. 1Shanghai Jiao Tong University  2EPIC Lab, SJTU 3Shanghai AI Laboratory  4Sun Yat-sen University zichen.wen@outlook.com, heconghui@pjlab.org.cn, zhanglinfeng@sjtu.edu.cn
@@ 1 Introduction
To solve this problem, abundant efforts have been made in token pruning (Chen et al., 2024; Zhang et al., 2024c; Liu et al., 2024d, 2025c, 2025a), which aims to reduce the number of input tokens in MLLMs. Usually, token pruning methods first introduce a carefully-designed criterion to measure the importance of a vision token, and then prune the redundant tokens, or merge the redundant tokens into fewer tokens. As a result, the following computation of the pruned tokens or the merged tokens can be removed or reduced, bringing efficiency in both computation and storage. For instance, some recent works show that more than 70% tokens can be pruned with a tolerant loss in accuracy (Chen et al., 2024). Most attractively, thanks to the natural ability of MLLMs to process tokens in different lengths, token pruning can be applied to most existing MLLMs with no need for additional training, and thus attracts great attention from both academic researchers and industrial developers.
@@
However, despite the popularity of token pruning, numerous foundational questions have long been overlooked and remain largely unexplored, giving rise to several surprising phenomena. For instance, Figure 1 demonstrates the comparison between two classical token pruning methods including FastV (Chen et al., 2024) and SparseVLM (Zhang et al., 2024c), and two naive baselines, including random token selection and direct average pooling on tokens. Surprisingly, the two baselines outperform the two well-designed token pruning methods in most benchmarks by a clear margin. This counterintuitive phenomenon may demonstrate that the current understanding of so-called important tokens is far away from the truth. Unfortunately, most recent works just focus on purchasing higher performance, while ignoring these questions, which may hinder the long-term development of token pruning.
@@ 2.1 Multimodal Large Language Models
The remarkable success of large language models (LLMs) (Radford et al., 2019; Brown et al., 2020; Wen et al., 2024) has spurred a growing trend of extending their advanced reasoning capabilities to multi-modal tasks, leading to the development of vision-language models (VLMs) (Huang et al., 2023; Driess et al., 2023; Liu et al., 2024c; Bai et al., 2023). These VLMs typically consist of a visual encoder (Radford et al., 2021) that serializes input image representations and an LLM responsible for text generation. To enable the LLM to process visual inputs, an alignment module is employed to bridge the gap between visual and textual modalities. This module can take various forms, such as an MLP projector or a more complex query-based network. While this integration allows the LLM to gain visual perception, it also introduces significant computational challenges due to the long sequences of visual tokens.
@@ 2.2 Visual Token Compression
Visual tokens are often significantly more numerous than text tokens, with higher spatial redundancy and lower information density. To address this issue, various methods have been proposed for reducing visual token counts in vision language models. For instance, some approaches modify model components, such as using context tokens in Q-Former (Li et al., 2024) or applying adaptive pooling at the patch level, but these typically require additional training and increase computational costs. Other techniques, like Token Merging (ToMe) (Bolya et al., 2023) and FastV (Chen et al., 2024), focus on reducing tokens without retraining by merging tokens or selecting important ones based on attention scores. SparseVLM Zhang et al. (2024c) incorporates text guidance through cross-modal attention to refine token selection. However, these methods often overlook hardware acceleration compatibility and fail to account for token duplication alongside token importance. Furthermore, while token pruning has been extensively explored in natural language processing and computer vision to improve inference efficiency, its application to VLMs remains under-explored. Existing pruning strategies, such as those in FastV and SparseVLM, rely on text-visual attention within large language models (LLMs) to evaluate token importance, which may not align well with actual visual token relevance.
@@ 3 Benchmarking
We begin by presenting the datasets, models, and pruning methods included in our study, along with the rationale behind the selection. Next, we outline the experimental setup and provide guidance on interpreting the results reported in our study. Finally, we analyze the findings, emphasizing notable patterns and offering insights that may inform future research in this area.
@@ 3.1 Models
We selected several representative MLLMs, including LLaVA-1.5-7B and 13B (Liu et al., 2024a), LLaVA-Next-7B (Liu et al., 2024b), and Qwen2-VL (Wang et al., 2024) series (7B-Instruct and 72B-Instruct). LLaVA-1.5-7B integrates CLIP and LLaMA for vision-language alignment via end-to-end training, employing MLP connectors to fuse visual-text features for multimodal reasoning. LLaVA-Next-7B enhances data efficiency and inference robustness with dynamic resolution and hierarchical feature integration, improving fine-grained visual understanding. Qwen2-VL series excel in high-resolution input processing and instruction-following, supporting complex tasks like document analysis and cross-modal in-context learning through unified vision-language representations.
@@ 3.3 Token Pruning Method
To rigorously evaluate the properties of visual token pruning, we select three representative and high-performing methods: FastV (Chen et al., 2024), SparseVLM Zhang et al. (2024c), and MustDrop (Liu et al., 2024d). FastV Chen et al. (2024) optimizes computational efficiency by learning adaptive attention patterns in early layers and pruning low-attention visual tokens post-layer 2 of LLMs, effectively reducing redundancy. SparseVLM (Zhang et al., 2024c) introduces a text-guided, training-free pruning mechanism that leverages self-attention matrices between text and visual tokens to assess importance. It maximizes sparsity while preserving semantically relevant tokens without additional parameters or fine-tuning. MustDrop (Liu et al., 2024d) addresses token redundancy across the entire model lifecycle. It merges spatially similar tokens during vision encoding, employs text-guided dual-attention filtering in prefilling, and implements output-aware KV cache compression during decoding. This multi-stage approach ensures balanced retention of critical tokens while enhancing inference efficiency. These methods exemplify diverse strategies for token pruning, spanning adaptive attention, text-guided sparsity, and lifecycle-aware optimization.
@@
[…]
Upper Bound, 576 Tokens  (100%)
Vanilla 63.3 68.9 62.3 1818 85.9 72.8 61.3 56.6 100%
[…]
Random 60.3 65.9 58.3 1767 80.6 71.4 54.3 57.6 95.5% (-4.5%)
Pooling 60.6 65.5 57.7 1742 83.6 71.3 56.3 56.6 95.8% (-4.2%)
Window FastV 59.5 65.5 57.7 1674 82.8 57.2 48.0 60.9 91.8% (-8.2%)
Vanilla FastV 57.7 53.9 46.8 1633 79.3 57.0 51.0 60.3 86.9% (-13.1%)
SparseVLM 57.9 63.8 55.8 1704 81.1 69.9 43.9 56.3 91.1% (-8.9%)
[…]
Random 57.5 62.6 54.4 1681 73.8 70.9 50.6 57.5 91.1% (-8.9%)
Pooling 55.4 58.3 51.4 1552 74.0 72.0 52.3 51.4 87.7% (-12.3%)
Window FastV 56.7 58.3 51.4 1599 76.5 55.4 43.2 59.5 85.7% (-14.3%)
Vanilla FastV 53.7 50.9 42.1 1567 69.3 56.8 47.1 59.2 81.6% (-18.4%)
SparseVLM 50.6 61.3 54.8 1402 65.0 69.0 22.7 54.5 79.7% (-20.3%)
Table 2: Performance Comparison of LLaVA-1.5-13B with Different Token Retention Strategies.
[…]
Upper Bound, 576 Tokens  (100%)
Vanilla 65.3 2521 87.4 91.6 82.8 100%
[…]
Random 63.9 2476 87.1 85.7 74.0 95.7% (-4.3%)
Pooling 63.1 2463 86.9 86.9 75.1 95.9% (-4.1%)
Window FastV 64.2 2445 88.5 85.9 75.4 96.3% (-3.7%)
Vanilla FastV 56.5 2219 80.9 85.3 75.6 90.3% (-9.7%)
[…]
Random 61.9 2394 85.5 79.3 64.5 90.4% (-9.6%)
Pooling 61.9 2391 84.3 81.1 65.9 90.9% (-9.1%)
Window FastV 61.9 2377 85.6 83.8 65.8 91.6% (-8.4%)
Vanilla FastV 55.7 2089 78.7 83.3 66.8 86.0% (-14.0%)
Table 3: Performance Comparison of Qwen2-VL-72B with Different Token Retention Strategies.
@@ 4 Token Pruning Revisited: Are Simple Methods Better?
When considering token pruning in multimodal large language models, two very basic methods naturally come to mind: random token pruning (hereafter referred to as Random) and token pooling (hereafter referred to as Pooling). Comparison with these two simple baselines is reliable evidence to demonstrate the significance of a well-designed token pruning method, yet has been ignored by most previous works. To address this gap, we investigated these two simple approaches in detail. Specifically, we conducted experiments on multiple widely-used benchmarks under pruning ratios of 75% and 87.5%, comparing Random and Pooling [^1] In the experiments, Pooling specifically refers to applying a pooling operation to the visual tokens at the second layer of the language model. Please refer to the specific implementation in Algorithm 4. with several recent token pruning methods (e.g., FastV and SparseVLM). As shown in Table 1, surprisingly, Random and Pooling outperformed carefully designed methods on nearly $2/3$ benchmarks. When scaling up to larger models, the experimental results of LLaVA-1.5-13B [^2] https://huggingface.co/liuhaotian/llava-v1.5-13b and Qwen2-VL-72B [^3] https://huggingface.co/Qwen/Qwen2-VL-72B-Instruct in Tables 2 and 3 also demonstrate the superior performance of the simple methods Random and Pooling. These surprising results shocked us since its inferior performance compared with random selection may demonstrate that we are on the wrong road toward the ideal token pruning methods.
@@ 4.1 Token Distribution: Spatial Uniformity Outperforms Position Bias
Figure 3: Sparse Visualization of Vanilla FastV and Window FastV with 25% Retained Visual Tokens.
@@ 5 Language in Visual Token Pruning: When and Why Does Language Matter?
Token pruning methods for multimodal models can be broadly categorized into two types: those guided by textual information (e.g., FastV Chen et al. (2024), SparseVLM Zhang et al. (2024c), MustDrop Liu et al. (2024d)) and those that rely solely on visual information (e.g., FasterVLM Zhang et al. (2024b)). While both approaches achieve comparable performance on common benchmarks, however, we hypothesize: Could it be that the importance of language information is not evident simply because there has been a lack of testing on tasks where language information is especially critical? To validate our hypothesis, we select a typical scenario: Visual Haystack.
@@ 5.1 Visual Token Pruning in Strongly Text-Guided Tasks
However, there are also recent works methods (Zhang et al., 2024b; Liu et al., 2025b) that perform pruning solely in ViT without textual information and reports better performance than FastV and SparseVLM in common VQA benchmarks.
@@ 6.1 Redundancy Criteria
Through the lens of mutual information (Latham and Roudi, 2009), we formulate this as maximizing information preservation between original tokens $\mathbf{X}$ and retained tokens $\mathbf{X^{\prime}}$:
$\max_{\mathcal{P}}I(\mathbf{X};\mathbf{X^{\prime}})=\mathcal{H}(\mathbf{X})-\mathcal{H}(\mathbf{X}|\mathbf{X^{\prime}}),$ (1)
where $\mathcal{P}$ denotes the pruning operator. This ensures the $\mathbf{X^{\prime}}$ retains maximal dependence on $\mathbf{X}$ under length constraint $\|\mathbf{X^{\prime}}\|=\|\mathbf{X}\|-\Delta L$. The formulation directly connects to the compression phase of the information bottleneck principle (Tishby et al., 2000), where $\mathcal{P}^{*}$ solves:
$\mathcal{P}^{*}=\operatorname*{arg\,min}_{\mathcal{P}}\|\mathbf{X^{\prime}}\|\quad\text{s.t.}\ I(\mathbf{X};\mathbf{X^{\prime}})\geq\gamma,$ (2)
@@ 6.2 Importance Criteria
In contrast, this task-oriented criterion explicitly considers the target output $\mathbf{Y}$. The goal shifts to preserving tokens critical for prediction accuracy, formalized through predictive sufficiency:
$I(\mathbf{X^{\prime}};\mathbf{Y})\geq I(\mathbf{X};\mathbf{Y})-\epsilon,$ (3)
@@
$\underbrace{I(\mathbf{X};\mathbf{Y})}_{\text{Original}}=\underbrace{I(\mathbf{X^{\prime}};\mathbf{Y})}_{\text{Pruned}}+\underbrace{I(\mathbf{X}\setminus\mathbf{X^{\prime}};\mathbf{Y}|\mathbf{X^{\prime}})}_{\text{Discarded}}.$ (4)
The bound $I(\mathbf{X}\setminus\mathbf{X^{\prime}};\mathbf{Y}|\mathbf{X^{\prime}})\leq\epsilon$ implies that discarded tokens provide negligible additional information about $\mathbf{Y}$ when conditioned on retained tokens. This captures the essence of importance - truly critical tokens contain non-decomposable predictive information.
@@
$\mathcal{R}(\beta)=\max_{\mathbf{X^{\prime}}}\left[I(\mathbf{X^{\prime}};\mathbf{Y})-\beta^{-1}I(\mathbf{X};\mathbf{X^{\prime}})\right],$ (5)
@@ 6.3 Empirical Validation of Adaptive Criteria Balancing
$\alpha\cdot\underbrace{I(x_{i};\mathbf{Y}|x_{\setminus i})}_{\text{Predictive Criticality}}+(1-\alpha)\cdot\underbrace{[1-I(x_{i};\mathbf{X}_{\setminus i})]}_{\text{Pattern Uniqueness}}.$ (6)
Here $I(x_{i};\mathbf{Y}|x_{\setminus i})$ measures a token’s unique predictive value, while $1-I(x_{i};\mathbf{X}_{\setminus i})$ quantifies its pattern distinctiveness.
@@
  • Perception-Dominant Tasks (MME, POPE) achieve peak performance at $\alpha=0.1$ and $0.0$, respectively, favoring redundancy-first pruning to maintain structural integrity ($\uparrow I(\mathbf{X};\mathbf{X^{\prime}})$).
@@ Appendix A Future Works
In this work, we have conducted an in-depth exploration of a series of issues related to token pruning. These include the position bias problem inherent in methods based on attention scores, the guiding role of linguistic information during token pruning, the importance and redundancy of tokens, as well as certain limitations in the evaluation of token pruning methods. Looking ahead, we plan to further expand the scope of our research by considering whether token pruning or token merging should be prioritized in the context of token reduction. Additionally, we aim to evaluate and analyze various token reduction methods on more challenging OCR benchmarks (Zhang et al., 2024a; Ouyang et al., 2024), particularly datasets featuring rich-text OCR images. This future work will not only deepen our understanding of token reduction strategies but also provide valuable insights into their practical applications in complex scenarios.